OpenStudy (anonymous):

can someone please help me, I have posted this before and could not get help..thanks Suppose that you have $10,000 to invest over a 3 year period. There are two accounts to choose from: 4% compounded monthly or 3.5% compounded continuously. a. Write the formula for the first account’s compound interest for n compoundings per year. b. Write the formula for the account with continuously compounded interest. c. Use the formulas and information above to solve for the balance of each investment after 3 years. d. Which investment account offers a better return on your money?

OpenStudy (anonymous):

\[1000(1+\frac{.04}{12})^{3\times 12}\] is the first one

OpenStudy (anonymous):

A=P(1+r/n)^nt A=Pe^rt plug in

OpenStudy (anonymous):

\[1000e^{.035\times 3}\] is the second

OpenStudy (anonymous):

now you need a calculator because you cannot do this without one. only suggestion is that when you use one compute \[3\times 12=36\] first so you just write \[1000(1+\frac{.04}{12})^{36}\] for the first one

OpenStudy (anonymous):

careful with parentheses on both of these